# A Survey on 3D Gaussian Splatting

### v2

@@ 5.1 Simultaneous Localization and Mapping (SLAM)
SLAM is a computational problem central to robotics and autonomous systems. It involves the challenge of a robot or a device understanding its position in an unknown environment while simultaneously mapping the environment’s layout [104]. SLAM is critical in various applications, including autonomous vehicles, augmented reality, and robotic navigation [105, 106]. The core of SLAM is to create a map of an unknown environment and determine the device’s location on this map in real-time. As a result, SLAM poses great challenges for computationally intensive scene representation techniques, yet serves as a good testbed for 3D GS.
3D GS enters the SLAM domain as an innovative approach for scene representation. Traditional SLAM systems often use point/surfel clouds or voxel grids to represent environments [107, 108, 109, 110, 111, 112, 113, 114]. In contrast, 3D GS utilizes anisotropic Gaussians to better represent environments. Recent innovative studies have employed 3D Gaussian Splatting in SLAM [115, 116, 117, 118, 119, 120, 121], showcasing the potential and versatility of this paradigm. An intuitive way is to use 3D Gaussian directly as the basic representation of the dense map and optimize the tracking process. For instance, GS-SLAM [118] adopted an adaptive strategy to add or remove 3D Gaussians so as to optimize scene geometry reconstruction and improve the mapping of previously observed areas. It also implemented a coarse-to-fine scheme for efficient camera pose tracking by selecting reliable 3D Gaussians. SplaTAM [119] integrated a straightforward online tracking and mapping method that leverages a silhouette mask for capturing the presence of scene density, facilitating fast rendering, dense optimization, and structured map expansion. GSSLAM [116] optimized 3D Gaussians directly for robust camera tracking and introduced geometric verification and regularization to tackle reconstruction ambiguities. On the other hand, designing advanced scene representations is worth exploring. For example, Photo-SLAM [115] proposed a hyper primitives map, which combines explicit geometric features for precise localization with implicit photometric features for texture mapping.
@@ 5.2 Dynamic Scene Reconstruction
To extend the concept of 3D GS to dynamic scenes, a straightforward approach is to incorporate the temporal dimension, allowing for the representation and rendering of scenes that change over time. 3D GS based methods [98, 97, 126, 127, 128, 129, 99, 130, 131, 132, 133, 134, 135, 136] for dynamic scene reconstruction can generally be divided into two main categories. The first category utilizes additional structural information like spatial MLPs or grids (Sec. 4.6). For example, Yang et al. [98] first proposed deformable 3D Gaussians tailored for dynamic scenes. These 3D Gaussians are learned in a canonical space and can be used to model spatial-temporal deformation with an implicit deformation field (implemented as an MLP). SC-GS [131] leveraged sparse control points and a deformation MLP to capture and represent the dynamics of 3D scenes. By decomposing motion into compact, learnable transformations applied to these control points, SC-GS achieved efficient and coherent motion modeling and allowed for motion editing. On the other hand, the second category is based on the idea that scene changes can be encoded into the 3D Gaussian representation with a specially designed rendering process (Sec. 4.5). For instance, Luiten et al. [130] introduced dynamic 3D Gaussians to model dynamic scenes by keeping the characteristics of 3D Gaussians (i.e., color, opacity, and size) constant over time while allowing their positions and orientations to change. Yang et al. [97] designed a 4D Gaussian representation, where additional properties are used to represent 4D rotations and spherindrical harmonics, to approximate the spatial-temporal volume of scenes.
@@ 5.3 AI-Generated Content (AIGC)
This explicit nature of 3D GS facilitates real-time rendering capabilities and unprecedented levels of control and editability, making it highly relevant for and AIGC applications. The explicit scene representations and differentiable rendering algorithms of 3D GS align perfectly with the requirements of AIGC for generating high-fidelity, real-time, and editable content, crucial for applications in virtual reality, interactive media, and beyond. Recent works have effectively utilized 3D GS in conjunction with fields like generative models [141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159], avatars [160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185], and scene editing [186, 187, 188, 95, 96, 94, 189, 131, 132, 133, 190, 191, 192]. For instance, DreamGaussian [142] accelerated the generation of photorealistic 3D assets from single-view image through a three-step process: a diffusion based generative GS process, followed by an efficient algorithm for mesh extraction from 3D Gaussians based on local density querying, and finally a UV-space refinement stage to improve texture details. By combining 3D GS with a parametric morphable face model, GaussianAvatars [169] offered enhanced fidelity and flexibility in avatar animation, significantly improving upon current methods in novel-view rendering and expression reenactment. GaussianEditor [186] introduced Gaussian semantic tracing for precise editing, hierarchical GS for stabilized results under random generative guidance, and a targeted editing strategy for efficient object removal and integration.
@@
TABLE I: Quantitative localization results6.1) on Replica [193], in terms of absolute trajectory error (ATE, cm). denotes numbers taken from [117]. (The three best scores are marked in red, blue, and green, respectively. These notes also apply to the other tables.)
@@
GS-SLAM [118] [CVPR24] 0.48 0.53 0.33 0.52 0.41 0.59 0.46 0.70 0.50
5.4 Autonomous Driving
Autonomous vehicles need to perceive and interpret their surroundings to navigate safely. This involves reconstructing the driving scene in real-time, accurately identifying static and dynamic objects, and understanding their spatial relationships and movements [209, 210, 211]. In dynamic driving scenes, the environment is continuously changing due to moving objects like other vehicles, pedestrians, or animals [212]. Accurately reconstructing these scenes in real-time is crucial for safe navigation but is challenging due to the complexity and variability of the elements involved. In autonomous driving, 3D GS can be utilized to reconstruct a scene by blending data points (such as those obtained from sensors like LiDAR) into a cohesive and continuous representation. This is particularly useful for handling the varying densities of data points and ensuring a smooth and accurate reconstruction of both the static background and dynamic objects in a scene. To reconstruct complex 3D scenes from sparse sensor data, especially at high speeds and with moving objects, mainstream frameworks separated the urban/street scene into static and dynamic elements, where the dynamic elements are modeled using a composite dynamic Gaussian graph [213], point clouds combined with semantic logits [214], or physically constrained models [215].
5.5 Endoscopic Scene Reconstruction
Surgical 3D reconstruction represents a fundamental task in the landscape of robot-assisted minimally invasive surgery, aimed at enhancing intraoperative navigation, preoperative planning, and educational simulations through precise modeling of dynamic surgical scenes. Pioneering the integration of cutting-edge dynamic radiance fields into this domain, recent advancements have focused on surmounting the inherent challenges of single-viewpoint video reconstructions – such as occlusions by surgical instruments and sparse viewpoint diversity within the confined spaces of endoscopic exploration [216, 217, 218]. Despite the progress, the call for high fidelity in tissue deformability and topological variance remains, coupled with the pressing demand for faster rendering to bridge the utility in real-time applications [219, 220, 221]. This synthesis of immediacy and precision in reconstructing deformable tissues from endoscopic videos is essential in propelling robotic surgery towards reduced patient trauma and augmented reality applications, ultimately fostering a more intuitive surgical environment and nurturing the future of surgical automation and robotic proficiency.
Compared to typical dynamic scene reconstruction, endoscopic scene reconstruction faces unique challenges, e.g., sparse training data due to constrained camera movement in confined space, unobserved regions due to tool occlusion, and pronounced non-rigid deformation of tissues. Existing approaches mainly used additional depth guidance to infer the geometry of tissues. For instance, Endo-4DGS [219] introduced a temporal component to 3D GS, utilizing a lightweight MLP to model temporal deformations, thereby facilitating the reconstruction of dynamic surgical scenes. By generating pseudo-depth maps from monocular views via an off-the-shelf depth estimator, it overcame the limitations of stereo camera setups and improved the depth-guided reconstruction process. Similarly, EndoGaussian [220] introduced two new strategies: holistic Gaussian initialization, which uses depth estimation models to predict depth maps from binocular/monocular image sequences, facilitating dense Gaussian initialization through pixel re-projection and combination; and spatiotemporal Gaussian tracking, which models scene dynamics using a deformation field, consisting of an efficient encoding voxel and a lightweight deformation decoder, and thus enables effective Gaussian tracking with reduced training and rendering efforts. EndoGS [221] also integrated depth-guided supervision with spatial-temporal weight masks and surface-aligned regularization terms to enhance the quality and speed of 3D tissue rendering while addressing tool occlusion. By utilizing a streamlined Gaussian representation combined with differentiable rasterization, EndoGSLAM [222] provided precise camera tracking, detailed 3D tissue reconstruction, and real-time visualization to support surgeons during minimally invasive procedures.
5.6 Medical Image
In the realm of medical imaging, X-ray stands as a foundational technique that employs electromagnetic radiation to capture images of the body’s internal structure [223]. This method hinges on the differential absorption of X-rays by various tissues, rendering visible the contrasts between bones, organs, and other internal elements without necessitating surgical intervention. Computed tomography (CT), an advanced iteration of X-ray imaging, enhances this process through the acquisition of multiple X-ray images from diverse angles around the patient. These images are then used to model a 3D representation (e.g., volumes), offering detailed cross-sectional views akin to slicing through the body layer by layer [224, 225]. This capability significantly amplifies the precision of diagnostics, allowing for the effective identification, monitoring, and management of medical conditions.
3D GS can be used directly for X-ray novel view synthesis, with a minor modification of the Gaussian’s attributes, i.e., from spherical harmonics to other functions to represent the absorption of X-ray radiation. Current works can be categorized according to their formulation of the scanned objects: i) w/o Volumes. X-Gaussian [226] proposed two main innovations: a redesigned radiative Gaussian point cloud model that captures the isotropic nature of X-ray radiation, utilizing a radiation intensity response function in place of spherical harmonics to model radiation intensity independently of view direction; and an angle-pose cuboid uniform initialization strategy for efficient point cloud initialization, leveraging X-ray scanner parameters to reduce training time significantly. ii) with Volumes. To reduce the radiation dose of traditional CT scans, Li et al. [227] used the capability of 3D Gaussians for sparse-view CT reconstruction. They utilized prior information from images reconstructed by filtered back projection to initialize the Gaussian attributes and adopted adaptive density control (Sec. 3.2.2) to optimize all 3D Gaussians in the projection space. Note that [227] used 3D Gaussians as the basis for the volumetric representation and does not involve the splatting process.
6 Performance Comparison
6.1 Performance Benchmarking: Localization
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve four recent papers which introduce 3D Gaussians into their systems [117, 116, 119, 118], as well as seven typical SLAM methods [194, 195, 196, 197, 198, 199, 200]
$\bullet$ Evaluation Metric: The root mean square error (RMSE) of the absolute trajectory error (ATE) is a commonly used metric in evaluating SLAM systems [228], which measures the root mean square of the Euclidean distances between the estimated and true positions over the entire trajectory of the robot or device.
$\bullet$ Result: As shown in Table I, the recent 3D Gaussians based localization algorithms have a clear advantage over existing NeRF based dense visual SLAM. For example, SplaTAM [119] achieves a trajectory error improvement of over 30%, decreasing it from 0.52cm to 0.36cm compared to the previous state-of-the-art (SOTA) [199]. We attribute this to the dense and accurate 3D Gaussians reconstructed for scenes, which can handle the noise of real sensors. This reveals that effective scene representations can improve the accuracy of localization tasks.
@@ 6.1 Performance Benchmarking: Localization
6.2 Performance Benchmarking: Rendering Quality in Static Scenes
$\bullet$ Result: Table II shows that 3D Gaussians based systems generally outperform the three dense SLAM competitors. For example, Gaussian-SLAM [117] establishes new SOTA and outperforms previous methods by a large margin. Compared to Point-SLAM [199], GSSLAM [116] is about 578 times faster in achieving very competitive accuracy. In contrast to previous method [199] that relies on depth information, such as depth-guided ray sampling, for synthesizing novel views, 3D GS based system [116] eliminates this need, allowing for efficient rendering for any given views with high fidelity.
6.3 Performance Benchmarking: Rendering Quality in Dynamic Scenes
$\bullet$ Dataset: D-NeRF [122] dataset includes videos with 50 to 200 frames each, captured from unique viewpoints. It features synthetic, animated objects in complex scenes, with non-Lambertian materials. The dataset provides 50 to 200 training images and 20 test images per scene, designed for evaluating models in monocular settings with varied camera poses.
$\bullet$ Result: From Table III we can observe that 3D GS based methods outperform existing SOTAs by a clear margin. The static version of 3D GS [12] fails to reconstruct dynamic scenes, resulting in a sharp drop in performance. By modeling the dynamics, CoGS [133] outperforms the SOTA method, FFDNeRF [125], by 5.22dB in terms of PSNR. These impressive results indicate the effectiveness of introducing additional properties or structured information to model the deformation of 3D Gaussians so as to model the scene dynamics.
@@ 6.3 Performance Benchmarking: Rendering Quality in Dynamic Scenes
6.4 Performance Benchmarking: Rendering Quality in Driving Scenes

### v3

@@ 5.1 Simultaneous Localization and Mapping (SLAM)
SLAM is a computational problem central to robotics and autonomous systems. It involves the challenge of a robot/device understanding its position in an unknown environment while simultaneously mapping the environment’s layout. SLAM is critical in various applications such as autonomous vehicles and robotic navigation. The core of SLAM is to create a map of an unknown environment and determine the device’s location on this map in real-time. As a result, SLAM poses great challenges for computationally intensive scene representation techniques, yet serves as a good testbed for 3D GS [156].
3D GS enters the SLAM domain as an innovative approach for scene representation. Traditional SLAM systems often use point/surfel clouds or voxel grids to represent environments [157, 158, 159, 160, 161, 162, 163, 164]. In contrast, 3D GS utilizes anisotropic Gaussians to better represent environments. Recent innovative studies have employed 3D GS in SLAM [108, 106, 107, 104, 105, 165, 110, 113], showcasing the potential and versatility of this paradigm. An intuitive way is to use 3D Gaussian as the basic representation of the dense map and optimize the tracking process. For instance, GS-SLAM [104] adopted an adaptive strategy to add or remove 3D Gaussians so as to optimize scene geometry reconstruction and improve the mapping of previously observed areas. Simiarly, Sun et al. [113] proposed additional regularization terms during the online mapping to avoid overfiting to the latest frame. SplaTAM [105] integrated a straightforward online tracking and mapping method that leverages a silhouette mask for capturing the presence of scene density, facilitating dense optimization and structured map expansion. On the other hand, designing advanced scene representations is worth exploring. For example, Photo-SLAM [108] proposed a hybrid Gaussian map, which combines explicit geometric features for precise localization with implicit photometric features for texture mapping. In addition, one interesting problem is how the Gaussian based map representations can help robotics tasks, and some early attempts in navigation have been made [166, 167, 168]. Though impressive, existing GS based SLAM systems still face persistent challenges such as dynamic elements, sensor noise, non-Lambertian objects, and depth ambiguities that require further exploration.
@@ 5.2 Dynamic Scene Reconstruction
To extend the concept of 3D GS to dynamic scenes, a straightforward approach is to incorporate the temporal dimension, allowing for the representation of scenes that change over time. 3D GS based methods [94, 93, 103, 173, 122, 174, 95, 118, 121, 120, 119, 175, 176, 123, 177] for dynamic scene reconstruction can generally be divided into two main categories. The first category utilizes additional structured information like spatial MLPs or grids (Sec. 4.6). For example, Yang et al. [94] first proposed deformable 3D Gaussians tailored for dynamic scenes. These 3D Gaussians are learned in a canonical space and can be used to model spatial-temporal deformation with an implicit deformation field (implemented as an MLP). GaGS [125] involves the voxelization of a set of Gaussian distributions, followed by the use of sparse convolutions to extract geometry-aware features, which are then utilized for deformation learning. On the other hand, the second category is based on the idea that scene changes can be encoded into the 3D Gaussian representation with a specially designed rendering process (Sec. 4.5). For instance, Luiten et al. [118] introduced dynamic 3D Gaussians to model dynamic scenes by keeping the properties of 3D Gaussians unchanged over time while allowing their positions and orientations to change. Yang et al. [93] designed a 4D Gaussian representation, where additional properties are used to represent 4D rotations and spherindrical harmonics, to approximate the spatial-temporal volume of scenes. While significant progress has been made in modeling dynamics and deformation at the Gaussian level, there is a compelling need to explore motion modeling at the object level, which might offer solutions to prevalent challenges like artifact reduction and the precise capture of fine-grained movements across long temporal sequences.
@@ 5.3 AI-Generated Content (AIGC)
This explicit nature of 3D GS facilitates real-time rendering capabilities and unprecedented levels of control and editability, making it highly relevant for and AIGC applications. The explicit scene representations and differentiable rendering algorithms of 3D GS align perfectly with the requirements of AIGC for generating high-fidelity, real-time, and editable content, crucial for applications in virtual reality, interactive media, and beyond. Recent works have effectively utilized 3D GS in conjunction with fields like generative models [182, 126, 127, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 128, 193, 194, 195, 129, 196, 197, 198, 199, 200, 201, 202, 203, 204], avatars [205, 206, 207, 130, 208, 209, 210, 211, 212, 213, 214, 131, 215, 132, 216, 217, 23, 218, 219, 220, 221, 222, 223, 224, 225, 226, 133, 227], and scene editing [134, 135, 228, 91, 92, 90, 229, 121, 120, 119, 230, 231, 232, 233, 234, 235, 236, 136, 237, 102]. For instance, DreamGaussian [126] accelerated the generation of photorealistic 3D assets from single-view image through a three-step process: a diffusion based generative GS process, followed by an efficient algorithm for mesh extraction from 3D Gaussians based on local density querying, and finally a UV-space refinement stage to improve texture details. By combining 3D GS with a parametric morphable face model, GaussianAvatars [213] offered enhanced fidelity and flexibility in avatar animation, significantly improving upon current methods in novel-view rendering and expression reenactment. To improve the efficacy of text instruction based editing, Chen et al. [134] devised a semantic tracker to track the editing target during training while Fang et al. [135] proposed to extract the region of interest corresponding to the instruction. These advancements hold promise for numerous industrial applications, including digital assetization (with mesh extraction), the generation of long-form videos (e.g., Sora), etc.
5.4 Autonomous Driving
Autonomous vehicles need to perceive and interpret their surroundings to navigate safely. This involves reconstructing the driving scene in real-time, accurately identifying static and dynamic objects, and understanding their spatial relationships and movements [246, 247, 248]. In dynamic driving scenes, the environment is continuously changing due to moving objects like other vehicles, pedestrians, or animals [249]. Accurately reconstructing these scenes in real-time is crucial for safe navigation but is challenging due to the complexity and variability of the elements involved. In autonomous driving, 3D GS can be utilized to reconstruct a scene by blending data points (such as those obtained from sensors like LiDAR) into a cohesive and continuous representation. This is particularly useful for handling the varying densities of data points and ensuring a smooth and accurate reconstruction of both the static background and dynamic objects in a scene. To reconstruct complex 3D scenes from sparse sensor data, especially at high speeds and with moving objects, mainstream frameworks separated the urban/street scene into static and dynamic elements, where the dynamic elements are modeled using a composite dynamic Gaussian graph [137], point clouds combined with semantic logits [138], or physically constrained models [139]. By delving deeper into physics- and semantics-aware 3D GS (see “physics- and semantics-aware scene representation” in Sec. 7), one can expect that 3D GS could ideally serve as the cornerstone for environment perception in autonomous driving.
5.5 Endoscopic Scene Reconstruction
Surgical 3D reconstruction represents a fundamental task in robot-assisted minimally invasive surgery, aimed at enhancing intraoperative navigation, preoperative planning, and educational simulations through precise modeling of dynamic surgical scenes. Pioneering the integration of cutting-edge dynamic radiance fields into this domain, recent advancements have focused on surmounting the inherent challenges of single-viewpoint video reconstructions – such as occlusions by surgical instruments and sparse viewpoint diversity within the confined spaces of endoscopic exploration [250, 251, 252]. Despite the progress, the call for high fidelity in tissue deformability and topological variance remains, coupled with the pressing demand for faster rendering to bridge the utility in applications sensitive to latency [140, 141, 142]. This synthesis of immediacy and precision in reconstructing deformable tissues from endoscopic videos is essential in propelling robotic surgery towards reduced patient trauma and AR/VR applications, ultimately fostering a more intuitive surgical environment and nurturing the future of surgical automation and robotic proficiency.
Compared to typical dynamic scene reconstruction, endoscopic scene reconstruction faces unique challenges, e.g., sparse training data, due to constrained camera movement in confined space, unobserved regions due to tool occlusion, and pronounced non-rigid deformation of tissues. Existing approaches mainly used additional depth guidance to infer the geometry of tissues [140, 141, 142]. For instance, EndoGS [142] integrated depth-guided supervision with spatial-temporal weight masks and surface-aligned regularization terms to enhance the quality and speed of 3D tissue rendering while addressing tool occlusion. EndoGaussian [141] introduced two new strategies: holistic Gaussian initialization for dense initialization and spatiotemporal Gaussian tracking for modeling surface dynamics. Zhao et al. [143] argued that these methods suffer from under-reconstruction and proposed to alleviate this problem from frequency perspectives. In addition, EndoGSLAM [116] and Gaussian Pancake [144] devised SLAM systems for endoscopic scenes and showed significant speed advantages. Note that present efforts are mainly directed towards reconstructing from a single viewpoint, which continues to face obstacles in surgical applications.
5.6 Large-scale Scene Reconstruction
Large-scale scene reconstruction is a critical component in fields such as autonomous driving, aerial surveying, and AR/VR, demanding both photorealistic visual quality and real-time rendering capabilities. Before the emergence of 3D GS, the task has been approached using NeRF based methods, which, while effective for smaller scenes, often fall short in detail and rendering speed when scaled to larger areas (e.g., over 1.5 $km^{2}$). Though 3D GS has demonstrated considerable advantages over NeRFs, the direct application of 3D GS to large-scale environments introduces significant challenges. 3D GS requires an immense number of Gaussians to maintain visual quality over extensive areas, leading to prohibitive GPU memory demands and considerable computational burdens during rendering. For instance, a scene spanning 2.7 $km^{2}$ may require over 20 million Gaussians, pushing the limits of even the most advanced hardware (e.g., NVIDIA A100 with 40GB memory) [146].
To address the highlighted challenges, researchers have made significant strides in two key areas: i) For training (learning 3D Gaussians for a scene), a divide-and-conquer strategy [146, 147, 148, 145] has been adopted, which segments a large scene into multiple, independent cells. This facilitates parallel optimization for expansive environments. An additional challenge lies in maintaining visual quality, as large-scale scenes often feature texture-less surfaces that can hamper the effectiveness of optimization such as Gaussian initialization and density control (Sec. 3.2). Enhancing the optimization algorithm presents a viable solution to mitigate this issue [45, 147]. ii) Regarding rendering, the adoption of the Level of Details (LoD) technique from computer graphics has proven instrumental. This technique adjusts the complexity of 3D scenes to balance visual quality with computational efficiency. Current implementations involve feeding only the essential Gaussians to the rasterizer [147], or designing explicit LoD structures like the Octree [148] and hierarchy [148]. Furthermore, integrating extra input modalities like LiDAR can further enhanced the reconstruction process [149, 150, 151]. One prominent challenge in large-scale scene reconstruction is limited capture, which could be alleviated by utilizing good priors (see “learning physical priors from large-scale data.” in Sec. 7).
6 Performance Comparison
@@
TABLE I: Quantitative localization results6.1) on Replica [253], in terms of absolute trajectory error (ATE, cm). (The three best scores are marked in red, blue, and green, respectively. These notes also apply to the other tables.)
@@
HFSLAM [113] [IROS24] 0.19 0.34 0.16 0.21 0.26 0.23 0.21 0.38 0.25
@@
6.1 Performance Benchmarking: Localization
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve five recent 3D GS based algorithms [107, 106, 105, 104, 113] and six typical SLAM methods [254, 255, 256, 257, 258, 259]
$\bullet$ Evaluation Metric: The root mean square error (RMSE) of the absolute trajectory error (ATE) is a commonly used metric in evaluating SLAM systems [260], which measures the root mean square of the Euclidean distances between the estimated and true positions over the entire trajectory.
$\bullet$ Result: As shown in Table I, the recent 3D Gaussians based localization algorithms have a clear advantage over existing NeRF based dense visual SLAM. For example, HFSLAM [113] achieves a trajectory error improvement of $\sim$50%, decreasing it from 0.52cm to 0.25cm compared to the previous state-of-the-art (SOTA) [258]. We attribute this to the dense and accurate 3D Gaussians reconstructed for scenes, which can handle the noise of real sensors. This reveals that effective scene representations can improve the accuracy of localization tasks.
6.2 Performance Benchmarking: Static Scenes
$\bullet$ Result: Table II shows that 3D Gaussians based systems generally outperform the three dense SLAM competitors. For example, Gaussian-SLAM [107] establishes new SOTA and outperforms previous methods by a large margin. Compared to Point-SLAM [258], GSSLAM [106] is about 578 times faster in achieving very competitive accuracy. In contrast to previous method [258] that relies on depth information, such as depth-guided ray sampling, for synthesizing novel views, 3D GS based system eliminates this need, allowing for high-fidelity rendering for any views.
6.3 Performance Benchmarking: Dynamic Scenes
$\bullet$ Dataset: D-NeRF [169] dataset includes videos with 50 to 200 frames each, captured from unique viewpoints. It features synthetic, animated objects in complex scenes, with non-Lambertian materials. The dataset provides 50 to 200 training images and 20 test images per scene, designed for evaluating models in the monocular setting.
$\bullet$ Result: From Table III we can observe that 3D GS based methods outperform existing SOTAs by a clear margin. The static version of 3D GS [10] fails to reconstruct dynamic scenes, resulting in a sharp drop in performance. By modeling the dynamics, D-3DGS [94] outperforms the SOTA method, FFDNeRF [172], by 6.83dB in terms of PSNR. These results indicate the effectiveness of introducing additional properties or structured information to model the deformation of Gaussians so as to model the scene dynamics.
@@ 6.3 Performance Benchmarking: Dynamic Scenes
6.4 Performance Benchmarking: Driving Scenes